# Activation degree thresholds and expressiveness of polynomial neural networks

### v3

@@ Introduction
While polynomial neural networks with fixed activation degree are not universal approximators [21], they become so when the activation degree is allowed to vary. An alternative measure of a polynomial neural network’s expressiveness is the dimension of its neurovariety, which can be defined as the maximum number of independent vectors in the tangent space at a generic point. From the perspective of algebraic geometry, the dimension of an algebraic variety is perhaps its most basic invariant. Characterizations of the neurovariety associated with a neural network represent a distinct description of a network’s expressiveness from universal approximation theorems. The dimension of the neurovariety is a simple and precise measure of the degrees of freedom of the model that admits exact computation. In contrast to universal approximation theorems, the dimension of the neurovariety is a measure of the expressiveness of a given polynomial neural network. Thus, studying polynomial network expressiveness through the dimension of the neurovariety complements universal approximation results. For an explicit toy example illustrating neurovariety dimension, see [30, Section A].
@@
The expressiveness of artificial neural networks of different structures and activation has been studied from several perspectives [31, 32, 37]. Substantial literature has shown the expressive power of deep networks, from the perspective of number of linear regions [32], universal approximation and VC dimensions [41, 37], a measure of complexity called “trajectory length” [31], and the exact class of functions representable by ReLU networks of various depths [20]. In another direction, [35] studied the effect of width of convolutional neural networks on the loss landscape, and [29, 19] provided bounds on network width for universal approximation for ReLU networks, while [24] provided such a bound for arbitrary activation functions. The expressiveness of neural networks has also been studied via their topological complexity, for example [7, 18]. In particular, [36] showed that networks with non-increasing architecture result in a connected decision region. ReLU expressiveness results have also been studied using algebraic geometry, specifically tropical geometry [9, 18, 20].
@@
This paper is structured as follows. In Section 1, we review key definitions for polynomial neural networks and introduce the notion of activation degree thresholds (Definition 10). In Section 2, we prove the existence of activation degree thresholds and determine the fibers of the parameter map of polynomial neural networks with high activation degree, resolving [25, Conjecture 16] and [28, Conjecture 5.2], raised by [14]. In Section 3, we prove the expressiveness of networks with equi-width architecture by computing the activation degree threshold— in particular, we remove any hypothesis involving “sufficiently high learning degree” by showing the activation degree threshold of these networks is one. In Section 4 we provide an outlook for the future, focusing on research directions opened by activation degree thresholds.
@@ 1.1 Background
$\begin{split}\Psi_{\mathbf{d},r}:\mathbb{R}^{d_{1}\times d_{0}}\times&\cdots\times\mathbb{R}^{d_{L}\times d_{L-1}}\to(\text{Sym}_{r^{L-1}}(\mathbb{R}^{d_{0}}))^{d_{L}},\\ \mathbf{w}&\mapsto p_{\mathbf{w}}=\begin{pmatrix}p_{\mathbf{w}_{1}}\\ \vdots\\ p_{\mathbf{w}_{d_{L}}}\end{pmatrix}.\end{split}$ (3)
@@
For all network architectures there exists a symmetry in the weight matrices, known as multi-homogeneity. It follows that the dimension of a fiber of the parameter map $\Psi_{\mathbf{d},r}$ is at least $\sum_{i=1}^{L-1}d_{i}$ ([25], Lemma 13).
@@
For all invertible diagonal matrices $D_{i}\in\mathbb{R}^{d_{i}\times d_{i}}$ and permutation matrices $P_{i}\in\mathbb{R}^{d_{i}\times d_{i}}$ ($i=1,\dots,L-1$), the parameter map $\Psi_{\mathbf{d},r}$ returns the same neural network under the replacement
@@
$W_{L}$ $\leftarrow W_{L}D_{L-1}^{-r}P_{L-1}^{T}.$
where $T$ denotes the matrix transpose. Thus the dimension of a generic pre-image of $\Psi_{\mathbf{d},r}$ is at least $\sum_{i=1}^{L-1}d_{i}$.
@@
$\mathrm{edim}\mathcal{V}_{\mathbf{d},r}:=\min$$\Bigg{\{}d_{L}+\sum_{i=0}^{L-1}(d_{i}d_{i+1}-d_{i+1}),\quad d_{L}\binom{d_{0}+r^{L-1}-1}{r^{L-1}}\Bigg{\}}.$ (4)
@@
However, the expected dimension is not always equal to the dimension, as demonstrated by the following example.
@@
The dimension of a neurovariety is subtle even in simple settings. For instance, when the architecture is $(d_{0},d_{1},1)$, a reinterpretation [25, Theorem 9] of the Alexander–Hirschowitz Theorem [2] determines which neurovarieties have the expected dimension. In [28], the authors coin the term defect for the difference between the expected dimension of a neurovariety and its dimension:
$\operatorname{defect}(\mathcal{V}_{\mathbf{d},r}):=\operatorname{edim}(\mathcal{V}_{\mathbf{d},r})-\dim(\mathcal{V}_{\mathbf{d},r}).$
Examples of neurovarieties $\mathcal{V}_{\mathbf{d},r}$ with positive defect are found in [28, Table 1] and [25].
@@
The notion of neurovariety dimension as a measure of network expressiveness has been utilized for designing neural networks. For example, [22] employs a hybrid network with monomial and LeakyReLU activation functions to determine the entanglement type of quantum states. Because their architecture is shallow, they apply the Alexander-Hirschowitz Theorem when choosing the number of neurons in each layer to guarantee that the associated neurovariety achieves the expected dimension. Such applications motivate the project of describing the neurovariety dimension for broader classes of architectures. Our Theorem 18 and Theorem 22 are examples of this kind of result.
@@ 1.2 Main result and example
Given an architecture $\mathbf{d}$, does there exist a non-negative integer $\tilde{r}$ such that the following holds:
$\dim\mathcal{V}_{\mathbf{d},r}=\mathrm{edim}\mathcal{V}_{\mathbf{d},r}\quad\text{for all}\quad r>\tilde{r}\,?$ (5)
When such a non-negative integer $\tilde{r}$ exists for (5) we have the following definition.
@@
The activation degree threshold of an architecture $\mathbf{d}=(d_{0},d_{1},\dots,d_{L})$ with $L>1$, is
$\operatorname{ActThr}(\mathbf{d}):=\min\{\tilde{r}\in\mathbb{N}_{\geq 0}:\dim\mathcal{V}_{\mathbf{d},r}=\mathrm{edim}\mathcal{V}_{\mathbf{d},r}\text{ for all }r>\tilde{r}\}.$
In other words, $\operatorname{ActThr}(\mathbf{d})$ is the smallest non-negative integer $\tilde{r}$ such that (5) holds.
With this terminology in hand, we state our main result, which shows the activation degree threshold is well defined when $\mathbf{d}$ has no entries equal to one.
@@
$\dim\mathcal{V}_{\mathbf{d},r}=\mathrm{edim}\mathcal{V}_{\mathbf{d},r}=d_{L}+\sum_{i=0}^{L-1}(d_{i}-1)d_{i+1}.$
@@
This example aims to illustrate the activation degree threshold by mapping the function space of neural networks with architecture $(3,2,1)$ and activation function $x^{r}$ for $r\in\mathbb{R}_{>0}$. To ensure these functions are well defined, we restrict $\mathbb{R}^{d_{0}}$ to the nonnegative orthant and assume that each weight matrix has nonnegative entries. Denote the corresponding parameter space and functional space as $\Theta^{+}$ and $\operatorname{Fun}^{+}_{r}\left(\mathbb{R}^{3},\mathbb{R}^{1}\right)$ respectively.
Note that each point $\mathbf{x}$ in $\mathbb{R}^{3}$ induces the map $E_{\mathbf{x}}:\operatorname{Fun}_{r}^{+}\left(\mathbb{R}^{3},\mathbb{R}^{1}\right)\to\mathbb{R}^{1}$ by evaluation $\psi\mapsto\psi(\mathbf{x})$. The neuromanifold $\operatorname{Fun}^{+}_{r}\left(\mathbb{R}^{3},\mathbb{R}^{1}\right)$ is known to have dimension at most $6$ for all $r$. Therefore, consider the six points $\mathbf{x}_{1},\dots,\mathbf{x}_{6}\in\mathbb{R}^{d_{0}}$ so that $(E_{\mathbf{x}_{1}},\dots,E_{\mathbf{x}_{6}})$ induces a linear map from $\operatorname{Fun}_{r}^{+}(\mathbb{R}_{\geq 0}^{3},\mathbb{R}_{\geq 0})$ to $\mathbb{R}_{\geq 0}^{6}$. Precomposing with the parameter map (1) we get a family of functions depending on $r$
$\mathbb{R}^{3\times 2}\times\mathbb{R}^{2\times 1}{\supset\Theta^{+}}\to{E\left(\operatorname{Fun}_{r}^{+}(\mathbb{R}_{\geq 0}^{3},\mathbb{R}_{\geq 0})\right)}\subset\mathbb{R}^{6},\quad\theta=(W_{1},W_{2})\mapsto E\left(F_{\theta}\right).$ (7)
The benefit of this setup is that it allows for a straightforward rank computation of the Jacobian matrix. For example, we calculate the Jacobian matrix of the map (7) with the following choices. The six points we choose to define $E$ are $(1,0,0)$, $(0,1,0)$, $(0,0,1)$, $(1,1,0)$, $(1,0,1)$, $(0,1,1)$, and we calculate the Jacobian of (7) at the point
$\left(\left[\begin{array}[]{cc}1&1\\ 1&2\\ 1&3\end{array}\right],\left[\begin{array}[]{c}1\\ 1\end{array}\right]\right)\in\mathbb{R}^{3\times 2}\times\mathbb{R}^{2\times 1}.$
@@
[$r&0&0&r&0&0&1&1\\ 0&r&0&0&r2^{r-1}&0&1&2^r\\ 0&0&r&0&0&r3^{r-1}&1&3^r\\ r2^{r-1}&r2^{r-1}&0&r3^{r-1}&r3^{r-1}&0&2^r&3^r\\ r2^{r-1}&0&r2^{r-1}&r4^{r-1}&0&r4^{r-1}&2^r&4^r\\ 0&r2^{r-1}&r2^{r-1}&0&r5^{r-1}&r5^{r-1}&2^r&5^r\\ \end{array}\right].$ (8)
Denote the $6\times 6$ submatrix on the left by $J(r)$. The positive integers where the $J(r)$ drops rank are $r=1,2$. Therefore the activation degree threshold of this architecture is less than or equal to three.
We visualize the functions $\det(J(r)$ and $\log|\det(J(r))|$ in Figure 1. Explicitly, the value of $\det(J(r))$ is the following polynomial exponential:
@@
Figure 1: The blue curve is the graph of $r\mapsto\frac{1}{r^{6}}\det(J(r))$ where $J(r)$ is the $6\times 6$ submatrix of (8). The red curve is $(r,\log|y|)$ where $(r,y)$ is on the blue curve. The orange dashed lines are $r=1$ and $r=2$.
@@ 2 Expressiveness for high activation degree
In this section we develop results on the linear independence of powers of polynomials (Section 2.1 and Section 2.2) to prove our main result in Section 2.3. In particular, our Theorem 18 completes the results in Corollary 17.
2.1 A prior number theoretic result by Newman-Slater
In this section, we recover a number theoretic result by Newman-Slater in [34]. The proof here slightly improves their bound.
@@ 2.1 A prior number theoretic result by Newman-Slater
A proof of this statement follows from modifications of the proof of [34, Section 4, Theorem]. We include it here for the benefit of the reader.
Consider the $k\times k$ Wronskian matrix $W:=W(R_{1}^{n},\dots,R_{k}^{n})$ where the $(i,j)$-th entry is the $(i-1)$-st derivative of $R_{j}^{n}$, which we denote by $\partial^{i-1}R_{j}^{n}$. With this setup, we have the following vector equation, with the first standard basis vector on the right hand side and $\mathbf{1}$ denoting the vector of ones:
@@
By letting $D$ be the $k\times k$ diagonal matrix with $(j,j)$ entry ${R_{j}^{-n+k-1}}$ we have
@@
$W\cdot D=\begin{bmatrix}R_{1}^{k-1}&\dots&R_{k}^{k-1}\\ \vdots&\dots&\vdots\\ \dfrac{\partial^{k-1}R_{1}^{n}}{R_{1}^{n-k+1}}&\dots&\dfrac{\partial^{k-1}R_{k}^{n}}{R_{k}^{n-k+1}}\\ \end{bmatrix}.$
@@
Assume each $R_{j}$ is written in lowest terms. After reordering the $R_{j}$’s as necessary, suppose $R_{1}$ has the largest degree numerator or denominator and denote this degree by $A$. Hence, the sum of the degrees of the numerator and denominator on the left hand side of (10) is between $A\cdot|n-k+1|$ and $2A\cdot|n-k+1|$.
@@
In other words, the sum of the degrees of the numerator and denominator, when taking the derivative once, increases by no more than $12\deg(g)-$ 1. As a consequence, the $(i,j)$-th entry of the matrix $W\cdot D$ when written in lowest terms is a rational function for which the sum of the degrees of the numerator and denominator is no more than $2A(k-1)+(2A-1)(i-1)$.
By the Leibniz formula for determinant, the sum of the degrees of the numerator and denominator for $\det(B)$ is no more than
$\sum_{i=2}^{k}\big{(}2A\cdot(k-1)+(2A-1)(i-1)\big{)}=$ $2\displaystyle 2A\cdot(k-1)^{2}\,+\,(2A-1)k(k-1)/$ 2
@@
$\sum_{i=1}^{k}\big{(}2A\cdot(k-1)+(2A-1)(i-1)\big{)}=$ $2\displaystyle 2A\cdot k(k-1)+(2A-1)k(k-1)/$ 2
$=$ $A\cdot(k-1)(3k)-k(k-1)/2.$
So we have
${\color[rgb]{0,0,0}A\cdot(n-k{+1})\leq A\cdot|n-k+1|}\leq\Big{(}A\cdot(k-1)(3k-2)-k(k-1)/2\Big{)}+\Big{(}A\cdot(k-1)(3k)-k(k-1)/2\Big{)}.$
@@
We have shown the inequality (11) holds when $R_{1}^{n},\dots,R_{k}^{n}$ are linearly independent. If these functions are linearly dependent, then we may find a subset $S$ of $\{1,\dots,k\}$ with cardinality $\ell$ less than $k$ such that $\sum_{j\in S}R_{j}^{n}=1$ and $\{R_{j}^{n}:j\in S\}$ is linearly independent. Applying the Wronskian argument in this setting proves
@@
Since $1<\ell<k$ implies $16{\ell}^{2}-6\ell+1<6{k}^{2}-6k+$ 1, the theorem holds.

### v4

@@ Introduction
While polynomial neural networks with fixed activation degree are not universal approximators [20], they become so when the activation degree is allowed to vary. An alternative measure of a polynomial neural network’s expressiveness is the dimension of its neurovariety, which can be defined as the maximum number of independent vectors in the tangent space at a generic point. From the perspective of algebraic geometry, the dimension of an algebraic variety is perhaps its most basic invariant. Characterizations of the neurovariety associated with a neural network represent a distinct description of a network’s expressiveness from universal approximation theorems. The dimension of the neurovariety is a simple and precise measure of the degrees of freedom of the model that admits exact computation. In contrast to universal approximation theorems, the dimension of the neurovariety is a measure of the expressiveness of a given polynomial neural network. Thus, studying polynomial network expressiveness through the dimension of the neurovariety complements universal approximation results. For an explicit toy example illustrating neurovariety dimension, see [29, Section A].
@@
Substantial literature has shown the expressive power of deep networks from the perspective of the number of linear regions [31], universal approximation and VC dimensions [42, 37], a measure of complexity called “trajectory length” [30], and the exact class of functions representable by ReLU networks of various depths [19]. In another direction, [34] studied the effect of the width of convolutional neural networks on the loss landscape, and [28, 18] provided limits on the network width for universal approximation for ReLU networks, while [23] provided such a limit for arbitrary activation functions. The expressiveness of neural networks has also been studied via their topological complexity, for example by [7, 17]. In particular, [35] showed that networks with non-increasing architecture result in a connected decision region. ReLU expressiveness results have also been studied using algebraic geometry, specifically tropical geometry [9, 17, 19].
@@
This paper is structured as follows. In Section 1, we review key definitions for polynomial neural networks and introduce the notion of activation degree thresholds (Definition 10). In Section 2, we prove the existence of activation degree thresholds and determine the fibers of the parameter map of polynomial neural networks with high activation degree, resolving [24, Conjecture 16] and [27, Conjecture 5.2], raised by [13]. In Section 3, we prove the expressiveness of networks with equi-width architecture by computing the activation degree threshold— in particular, we remove any hypothesis involving “sufficiently high learning degree” by showing the activation degree threshold of these networks is one. In Section 4 we provide an outlook for the future, focusing on research directions opened by activation degree thresholds.
@@ 1.1 Background
$\begin{split}\Psi_{\mathbf{d},r}:\mathbb{R}^{d_{1}\times d_{0}}\times&\cdots\times\mathbb{R}^{d_{L}\times d_{L-1}}\to(\text{Sym}_{r^{L-1}}(\mathbb{R}^{d_{0}}))^{d_{L}},\\ \mathbf{w}&\mapsto p_{\mathbf{w}}.\end{split}$ (3)
@@
The neurovariety $\mathcal{V}_{\mathbf{d},r}$ is a real variety parameterized by the polynomial map $\Psi_{\mathbf{d},r}:\mathbb{R}^{N}\to\mathbb{R}^{M}$ where $N=\sum_{i=0}^{L-1}d_{i}d_{i+1}$ and $M=d_{L}\binom{r^{L-1}+d_{0}-1}{d_{0}-1}$. Extending the domain of $\Psi_{\mathbf{d},r}$ to $\mathbb{C}^{N}$ results in the map
$\Psi_{{\mathbb{C},\mathbf{d},r}}:\mathbb{C}^{N}\to\mathbb{C}^{M}.$ (4)
Denote the Zariski closure of the image of $\Psi_{{\mathbb{C},\mathbf{d},r}}$ by $Y$. By [41, Theorem 12.2], the dimension of the complex variety $Y$ and the dimension of the real variety $\mathcal{V}_{\mathbf{d},r}$ agree if $Y$ contains a smooth real point. Since $Y$ is parameterized by polynomials with real coefficients, it follows that $Y$ contains a smooth real point and
$\operatorname{dim}{Y}=\operatorname{dim}{\mathcal{V}_{\mathbf{d},r}}.$ (5)
Equation (5) puts complex algebraic geometry at the disposal of statements concerning the dimension of $\mathcal{V}_{\mathbf{d},r}$. We exploit this in Theorems 18 and 22. Our approach relies on understanding the dimension of the fiber of $\Psi_{{\mathbb{C},\mathbf{d},r}}$ over $y\in\Psi_{{\mathbb{C},\mathbf{d},r}}(\mathbb{C}^{N})$:
$\Psi_{{\mathbb{C},\mathbf{d},r}}^{-1}(y):=\{w\in\mathbb{C}^{N}:\Psi_{{\mathbb{C},\mathbf{d},r}}(w)=y\}.$
For all network architectures there exists a symmetry in the weight matrices, known as multi-homogeneity.
@@
For all invertible diagonal matrices $D_{i}\in\mathbb{R}^{d_{i}\times d_{i}}$ and permutation matrices $P_{i}\in\mathbb{Z}^{d_{i}\times d_{i}}$ ($i=1,\dots,L-1$), the parameter map $\Psi_{\mathbf{d},r}$ returns the same neural network under the replacement
@@
$W_{L}$ $\leftarrow W_{L}D_{L-1}^{-r}P_{L-1}^{T},$
where $T$ denotes the matrix transpose.
By letting the diagonal matrices $D_{i}$ have complex entries, the multi-homogeneity lemma has known consequences for the fibers of $\Psi_{{\mathbb{C},\mathbf{d},r}}$. Namely, the fiber of $\Psi_{{\mathbb{C},\mathbf{d},r}}$ over $y=\Psi_{{\mathbb{C},\mathbf{d},r}}(W_{1},\dots,W_{L})$ satisfies this containment:
[…]
and the dimension of a general fiber is at least $\sum_{i=1}^{L-1}d_{i}$ [24, Lemma 13]. For clarity, we remark that when $L=1$, the right-hand side of (6) simplifies to $\{W_{1}\}$, and the containment becomes an equality. We are careful to state when a polynomial map is over a real or complex domain to avoid ambiguity when discussing the dimension of preimages that are not general.
@@
$\operatorname{edim}{\mathcal{V}_{\mathbf{d},r}}:=\min$ $\Bigg{\{}d_{L}+\sum_{i=0}^{L-1}(d_{i}d_{i+1}-d_{i+1}),\quad d_{L}\binom{r^{L-1}+d_{0}-1}{d_{0}-1}\Bigg{\}}$. (7)
@@
The expected dimension is not always equal to the dimension, as the following example demonstrates.
@@
The dimension of a neurovariety is subtle even in simple settings. For instance, when the architecture is $(d_{0},d_{1},1)$, a reinterpretation [24, Theorem 9] of the Alexander–Hirschowitz Theorem [2] determines which neurovarieties have the expected dimension. In [27], the authors coin the term defect for the difference between the expected dimension of a neurovariety and its dimension:
$\operatorname{defect}(\mathcal{V}_{\mathbf{d},r}):=\operatorname{edim}{\mathcal{V}_{\mathbf{d},r}}-\operatorname{dim}{\mathcal{V}_{\mathbf{d},r}}.$
Examples of neurovarieties with positive defect are found in [27, Table 1] and in [24] as well.
@@
The notion of neurovariety dimension as a measure of network expressiveness has been utilized for designing neural networks. For example, [21] employs a hybrid network with monomial and LeakyReLU activation functions to determine the entanglement type of quantum states. Because their architecture is shallow, they apply the Alexander–Hirschowitz Theorem when choosing the number of neurons in each layer to guarantee that the associated neurovariety achieves the expected dimension. Such applications motivate the project of describing the neurovariety dimension for broader classes of architectures. Our Theorem 18 and Theorem 22 are examples of this kind of result.
@@ 1.2 Main result and example
Given an architecture $\mathbf{d}$, does there exist a nonnegative integer $\tilde{r}$ such that the following holds:
$\operatorname{dim}{\mathcal{V}_{\mathbf{d},r}}=\operatorname{edim}{\mathcal{V}_{\mathbf{d},r}}\quad\text{for all}\quad r>\tilde{r}\,$? (8)
When such a nonnegative integer $\tilde{r}$ exists for (8) we have the following definition.
@@
The activation degree threshold of an architecture $\mathbf{d}=(d_{0},d_{1},\dots,d_{L})$ with $L>1$ is
$\operatorname{ActThr}(\mathbf{d}):=\min\{\tilde{r}\in\mathbb{N}_{\geq 0}:\operatorname{dim}{\mathcal{V}_{\mathbf{d},r}}=\operatorname{edim}{\mathcal{V}_{\mathbf{d},r}}\text{ for all }r>\tilde{r}\}$.
In other words, $\operatorname{ActThr}(\mathbf{d})$ is the smallest nonnegative integer $\tilde{r}$ such that (8) holds.
With this terminology in hand, we state our main result, which shows that the activation degree threshold is well-defined when $\mathbf{d}$ has no entries equal to one.
@@
$\operatorname{dim}{\mathcal{V}_{\mathbf{d},r}}=\operatorname{edim}{\mathcal{V}_{\mathbf{d},r}}=d_{L}+\sum_{i=0}^{L-1}(d_{i}-1)d_{i+1}.$
@@
This example aims to illustrate the activation degree threshold by mapping the function space of neural networks with architecture $\mathbf{d}=(3,2,1)$ and activation function $x^{r}$ for $r\in\mathbb{R}_{>0}$. To ensure these functions are well-defined, we restrict $\mathbb{R}^{d_{0}}$ to the nonnegative orthant and assume that each weight matrix has nonnegative entries.
Denote the corresponding parameter space as $\Theta^{+}$ and the neuromanifold as $\operatorname{Fun}^{+}_{r}\left(\mathbb{R}^{3}_{\geq 0},\mathbb{R}_{\geq 0}\right)$. The neuromanifold $\operatorname{Fun}^{+}_{r}\left(\mathbb{R}^{3}_{\geq 0},\mathbb{R}_{\geq 0}\right)$ has dimension at most six for all $r\in\mathbb{R}_{>0}$ because any element in it is of the form
$(ax_{1}+bx_{2}+cx_{3})^{r}+(a^{\prime}x_{1}+b^{\prime}x_{2}+c^{\prime}x_{3})^{r}.$
Note that each point $\mathbf{z}$ in $\mathbb{R}^{3}$ induces an evaluation map
$\operatorname{Fun}_{r}^{+}\left(\mathbb{R}^{3}_{\geq 0},\mathbb{R}_{\geq 0}\right)\to\mathbb{R}_{\geq 0},\quad F\mapsto F(\mathbf{z}).$
Therefore, consider six points $\mathbf{z}_{1},\dots,\mathbf{z}_{6}\in\mathbb{R}^{3}$ and define the function
$E:\operatorname{Fun}_{r}^{+}(\mathbb{R}_{\geq 0}^{3},\mathbb{R}_{\geq 0})\to\mathbb{R}_{\geq 0}^{6}$
by $F\mapsto(F({\mathbf{z}_{1}}),\dots,F({\mathbf{z}_{6}}))$. Precomposing with the parameter map (1), we get a family of functions depending on $r$:
$\mathbb{R}^{2\times 3}\times\mathbb{R}^{1\times 2}{\supset\Theta^{+}}\to E\left(\operatorname{Fun}_{r}^{+}(\mathbb{R}_{\geq 0}^{3},\mathbb{R}_{\geq 0})\right)\subset\mathbb{R}^{6},\quad{\boldsymbol{\theta}}=(W_{1},W_{2})\mapsto E\left(F_{\boldsymbol{\theta}}\right).$ (10)
The benefit of this setup is that it allows for a straightforward rank computation of the Jacobian matrix. For example, we calculate the Jacobian matrix of the map (10) with the following choices. The six points we choose to define $E$ are
(1,0,0),(0,1,0),(0,0,1),(1,1,0),(1,0,1),(0,1,1),(1,0,0),\,\,(0,1,0),\,\,(0,0,1),\,\,(1,1,0),\,\,(1,0,1),\,\,(0,1,1),
and we calculate the Jacobian of (10) at the point
$\left(\left[\begin{array}[]{ccc}1&1&1\\ 1&2&3\end{array}\right],\left[\begin{array}[]{cc}1&1\end{array}\right]\right)\in\mathbb{R}^{2\times 3}\times\mathbb{R}^{1\times 2}.$
@@
[…]
Denote the $6\times 6$ submatrix on the left by $J(r)$. The positive integers where the $J(r)$ drops rank are $r=1,2$. Therefore the activation degree threshold of this architecture is less than or equal to two. When $r=2$, $F_{\boldsymbol{\theta}}$ is the sum of two squares, hence a degenerate quadratic form in $3$ variables. Therefore it has at most $5$ degrees of freedom, and so the threshold is exactly two.
We visualize the functions $\det(J(r))$ and $\log|\det(J(r))|$ in Figure 1. Explicitly, the value of $\det(J(r))$ is the following polynomial exponential:
@@
Figure 1: The blue curve is the graph of $r\mapsto\det(J(r))$ where $J(r)$ is the $6\times 6$ submatrix of (11). The red curve is $(r,\log|y|)$ where $(r,y)$ is on the blue curve. The orange dashed lines are $r=1$ and $r=2$.
@@ 2 Expressiveness for high activation degree
In this section, we develop results on the linear independence of powers of polynomials (Section 2.1 and Section 2.2) to prove our main result in Section 2.3. In particular, our Theorem 18 completes the results in Corollary 17. Our Corollary 17 completes the proof of [24, Theorem 14].
2.1 A prior number theoretic result by Newman–Slater
In this section, we recover a number-theoretic result stated by Newman–Slater in [33]. The proof here slightly improves their bound (which is $n\leq 8k^{2}$ in our notation).
@@ 2.1 A prior number theoretic result by Newman–Slater
A proof of this statement follows from modifications of the proof of [33, Section 4, Theorem]. We include it here for the benefit of the reader.
Consider the $k\times k$ Wronskian matrix $W:=W(R_{1}^{n},\dots,R_{k}^{n})$ where the $(i,j)$-th entry is the (i1)(i-1)-st derivative of $R_{j}^{n}$, which we denote by $\operatorname{d}^{i-1}R_{j}^{n}$. With this setup, we have the following vector equation, with the first standard basis vector on the right-hand side and $\mathbf{1}$ denoting the vector of ones:
@@
By letting $D$ be the $k\times k$ diagonal matrix with $(j,j)$ entry $1/{R_{j}^{n+1-k}}$ we have
@@
$W\cdot D=\begin{bmatrix}\dfrac{1}{R_{1}^{1-k}}&\dots&\dfrac{1}{R_{k}^{1-k}}\\ \vdots&&\vdots\\ \dfrac{\operatorname{d}^{k-1}R_{1}^{n}}{R_{1}^{n+1-k}}&\dots&\dfrac{\operatorname{d}^{k-1}R_{k}^{n}}{R_{k}^{n+1-k}}\\ \end{bmatrix}$.
@@
Assume each $R_{j}$ is written in lowest terms. After reordering the $R_{j}$’s as necessary, suppose $R_{1}$ has the largest degree numerator or denominator and denote this degree by $A$. Hence, the sum of the degrees of the numerator and denominator on the left-hand side of (13) is between $A\cdot|n-k+1|$ and $2A\cdot|n-k+1|$.
@@
In other words, the sum of the degrees of the numerator and denominator, when taking the derivative once, increases by no more than $2\deg(g)-1$. As a consequence, the $(i,j)$-th entry of the matrix $W\cdot D$ when written in lowest terms is a rational function for which the sum of the degrees of the numerator and denominator is no more than 2A(k1)+(2A1)(i1)2A(k-1)+(2A-1)(i-1). By the Leibniz formula for the determinant, the sum of the degrees of the numerator and denominator for $\det(B)$ is no more than
$\sum_{i=2}^{k}\big{(}2A\cdot(k-1)+(2A-1)(i-1)\big{)}=$ $2A\cdot(k-1)^{2}\,+\,(2A-1)k(k-1)/2$
@@
$\sum_{i=1}^{k}\big{(}2A\cdot(k-1)+(2A-1)(i-1)\big{)}=$ $2A\cdot k(k-1)+(2A-1)k(k-1)/2$
$=$ $A\cdot(k-1)(3k)-k(k-1)/2.\$
Using (13), we have
$A\cdot(n-k{+1})$
$\hskip-36.135pt\leq A\cdot|n-k+1|$
$\hskip-36.135pt\leq\Big{(}A\cdot(k-1)(3k-2)-k(k-1)/2\Big{)}+\Big{(}A\cdot(k-1)(3k)-k(k-1)/2\Big{)}.$
@@
We have shown the inequality (14) holds when $R_{1}^{n},\dots,R_{k}^{n}$ are linearly independent. If these functions are linearly dependent, then we may find a subset $S$ of $\{1,\dots,k\}$ with cardinality $\ell$ less than $k$ such that $\sum_{j\in S}R_{j}^{n}=1$ (after potentially rescaling the $R_{j}^{n}$) and $\{R_{j}^{n}:j\in S\}$ is linearly independent. Applying the Wronskian argument in this setting proves
@@
Since $1<\ell<k$ implies $6{\ell}^{2}-6\ell+1<6{k}^{2}-6k+1$, the theorem holds. ∎